Mathematics 241 – Multivariable Calculus Quiz 3 – September 20, 2013

Your Name:

A) (10) Find α′(t) for the parametric curve α(t) = (\frac{2t}{t^{2}+1}, sin(3t))

B) (10) Give a parametrization of the tangent line to the curve from part A at the point α(0).